OpenStudy (kayla2001):

the total was $29.00 after adding a 4.00 TIP To the bill for a family of five, each of whom ordered a hot dog. what was the price of a hotdog

OpenStudy (jojo4eva):

the price of the hot dog then would have been 5 dollars

OpenStudy (jojo4eva):

5*5=25+4=29

OpenStudy (jojo4eva):

so each hot dog was indeed $5